Kenco purchased a new CNC lathe 6 years ago that had an installed cash cost of $475,000. The lathe was being depreciated over 8 years using straight line deprecation without taking salvage value into consideration. On 12/31 of the sixth year, Kenco sold the lathe to a machine tool dealer for $85,000.

            Required:       Assume that lathe has been depreciated over six years and

that accumulated depreciation is up to date. Make the

required journal entry to record the transaction in

Kenco’saccounting system.
Answer #1

Annual depreciation = (475000-0)/8 = $59375

Accumulated depreciation upto 6th year = 59375*6 = $356250

Accounts Debit Credit
Cash $85000
Accumulated depreciation - Lathe $356250
Loss on sale of Lathe $33750
Lathe $475000